OpenStudy (anonymous):

Is the equation y = x3 even, odd, or neither?

OpenStudy (wwe123):

it odd

OpenStudy (anonymous):

f(-x) = -f(x) means it's odd f(-x) = f(x) means it's even can you answer now ?

OpenStudy (anonymous):

its odd?

OpenStudy (anonymous):

nothings negative so its eve or neighter?

OpenStudy (anonymous):

*even

OpenStudy (anonymous):

lets do this f(-x) = (-x)^3 = - (x^3) = -f(x) its odd

OpenStudy (anonymous):

aight thanks